# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ 2.3. Non-Image Data
Due to the ubiquity of the raster image format in remotely sensed data, current RS foundation models tend to be focused on using images to train, but a variety of other modalities do exist. Text, geographic location, and digital elevation maps (DEM) are all among potential additional modalities available in remote sensing at scale, and it is conceivable that derivative information products such as population movement trajectories, social media data, and vector maps may be incorporated in enhancing foundation models. Any dataset with sufficient spatial coverage that contains geographic coordinates and a timestamp can feasibly be compared with corresponding satellite imagery in order to further improve a model’s ability to interpret how satellite imagery pertains to ground level information, resulting in a broad swathe of different modalities available for RS models to leverage. With an eye on these additional potential modalities, the rest of this manuscript focuses on the present and immediate future of foundation models in remote sensing with fundamental design and recommendations to advance the field.
@@ 3.1. Negative Sampling Models
One successful approach to SSL in CV is contrastive learning via negative sampling (Chen et al., 2020). Contrastive learning augments one image in two different ways, then a model is trained to recognize whether two images are augmentations of each other, i.e., are a positive pair. Any images that are not augmentations of one another are defined as negative pairs. For every batch, an image will have only one positive pair, and n-1 negative pairs, where n is the batch size. This approach can be thought of as a classification task where the model trains on classifying two augmentations as either a positive or negative pair. Within a given batch of samples, normalized-temperature cross entropy loss (NT-Xent loss) (Chen et al., 2020) pulls together embeddings from the same origin and pushes apart embeddings that came from different origins, resulting in an encoder that learns sophisticated embeddings of images without the need for supervised learning. The encoders then can be ported to other downstream tasks, while reducing the need for labeled samples.
@@
SimCLR (Chen et al., 2020) achieved great success with this approach; SimCLR trained on ImageNet ILSVRC-2012 (Russakovsky et al., 2015) and introduced a non-linear projection head to improve the embedding quality of the layer before it. They also discovered that augmentations like color jitter and random cropping were particularly effective in forcing the model to learn embeddings that represented the data as a whole, not just dominant features like color distribution. Momentum Contrast (MoCo) (He et al., 2020) also utilizes contrastive learning, but maintains a dynamic memory bank or queue of negative samples, and updates it using a momentum encoder. This approach allows MoCo to effectively utilize smaller batch sizes than SimCLR, reducing memory requirements. Contrastive Language Image Pre-training (CLIP) (Radford et al., 2021) is a contrastive learning framework that treats the text modality as a naturally occurring augmentation of a given image and trains two encoders, one for the text modality and one for the image modality. Given the multi-modal nature of this approach, contrastive learning to bring positive pairs together results in an alignment of the two modalities in the same embedding space. All of these models introduce concepts that have been leveraged by recent foundation models within the RS domain (Wang et al., 2022).
Two small diagrams that depict contrastive learning with negative sampling or distillation. In negative sampling, we see a batch of images augmented, passed through the same encoder, and then compared. In distillation, we see a single image augmented, passed through two different encoders, and then compared.
3.2. Distillation Networks
@@
While contrastive learning via negative sampling achieves SOTA performance in CV, it generally requires large batch sizes in order for the loss to effectively contrast against a large number of negative samples. The need for a large batch size results in significant hardware requirements, which is an obstacle to research for many teams. Distillation networks such as Bootstrap Your Own Latent (BYOL) (Grill et al., 2020) and Self-Distillation with No Labels (DINO) (Caron et al., 2021) were developed to address those hardware constraints.
Whereas negative sampling approaches can be thought of as a training on a classification task, distillation networks use a generative task as their objective. Distillation networks contain two encoders, dubbed the “teacher” and the “student”. Both are fed in augmented versions of an image. The teacher encoder outputs an embedding, and the student encoder is paired with a decoder that is responsible for generating a matching embedding. Then, generative loss is performed between the teacher and student embeddings and back propagated into the student network. The teacher network learns via the weighted average update system that was introduced in MoCo.
@@
The generative nature of the student network’s task forces the student network to learn effective representations without requiring negative samples. This method reduces the reliance on large batch sizes and extensive memory, making it more resource-friendly. It also outperforms comparable negative sampling models with fewer parameters (Grill et al., 2020). However, there is possibility of embedding collapse, as both the teacher and student networks learn based on the same loss (Chen and He, 2021). Embedding collapse describes a condition where a model generates embeddings that span a low-dimension subspace of the latent space, sometimes becoming so extreme as to only span a single point; this represents a significant loss of ability to capture information in embeddings (Jing et al., 2022). The asymmetry introduced by the weighted average update system for the teacher network prevents embedding collapse, but this phenomena is not well understood.
3.3. Masked Auto-Encoders
The Masked Auto-Encoder (MAE) approach is prevalent in NLP foundation models, such as BERT (Koroteev, 2021) and GPT (OpenAI et al., 2024). The notion is that by masking parts of a sample, running the masked sample through an encoder, then asking a decoder to reconstruct the original sample from the resulting embeddings, the encoder learns effective embeddings of the data. This approach has recently been extended to CV with models like MAE (He et al., 2021) achieving SOTA performance. As with contrastive learning approaches, MAE takes steps to ensure that learned embeddings are representative of the data as a whole. This takes the form of masking roughly 75% of the data to prevent trivial reconstruction from patch context (He et al., 2021). The generative nature of the task, like that of distillation networks, means that the batch size can be relatively small, making it relatively resource-friendly with moderate computing resources.
@@ 3.3. Masked Auto-Encoders
Masking approaches with established frameworks, such as Convolutional Neural Networks (CNNs) or Recursive Neural Networks (RNNs), replace a patch with a mask token, then feed in the whole image, including masked patches, into a network. But the usage of ViTs, which include positional embeddings as well as patch embeddings, allows for training pipelines to only feed encoders unmasked patches, shrinking the size of the necessary input layers and dramatically speeding up training time (He et al., 2021). Similarly, utilization of a lightweight decoder also forces the encoder to learn more accurate representations while speeding up runtime (He et al., 2021). These optimizations mean that the MAE approach scales remarkably well, resulting in models that can contain billions of parameters (Dehghani et al., 2023). Current CV research has identified a gap in the number of parameters between SOTA NLP and CV models and plans to improve the parameter count in CV foundation models even further (Dehghani et al., 2023).
A diagram depicting masked auto-encoding, where a given image is obscured by a large amount, and then passed through an encoder to generate an embedding. That embedding is then passed through a decoder which attempts to recreate the original image.
@@ 4. Negative Sampling Models in Remote Sensing
While the underlying principles of negative sampling generally remain the same from augmentations applied in CV to RS, the shared location of different imagery acquired at different times or by different sensors enables RS models to lean heavily on different sensors’ imagery or same sensor’s different temporal passes as naturally occurring augmentations of the same image (Jain et al., 2022a). This enables two main advantages: it lets a model leverage and fuse imagery that comes from different RS sensors or different temporal acquisitions that is available for the same locations, and it reduces the time required to do hyper-parameter tuning for augmentations. The latter often adds significant overhead to training contrastive learning approaches (Jain et al., 2022b).
4.1. Temporal Acquisitions as Augmentation
@@
Pre-trained on RGB bands from Sentinel-2, Seasonal Contrast (SeCo) (Mañas et al., 2021) borrows heavily from SimCLR’s general framework and outperforms CV models such as MoCo on downstream land cover classification tasks, indicating that there is implicit benefit in creating foundation models for the remote sensing domain as opposed to borrowing models directly from the computer vision field. Instead of creating hand-crafted image augmentations outlined in SimCLR (Chen et al., 2020), such as color jitter or gaussian blur, SeCo treats images of the same location at different points in time as natural augmentations of one another. However, SeCo does not simply rely on natural augmentations; it also utilizes random cropping, random flipping, and color jitter to prevent the model from learning trivial representations of the data. The resulting model is essentially season- or time-agnostic, meaning it provides a general representation of data regardless of acquisition-time-of-year. While this is helpful in identifying land cover or land use which does not change from season to season, it may not fit scenarios where retaining seasonal differences in embeddings for downstream tasks is essential, such as time series forecasting for projecting forest fire recovery in an area.
@@
SeCo’s approach of contrastive learning between spatially aligned, temporally different RS image pairs has also been explored in Geographically Aware Self Supervised Learning (GASSL) (Ayush et al., 2021). However, this approach trains on the Functional Map of the World (FMoW) (Christie et al., 2018) dataset, which contains multispectral imagery from the DigitalGlobe constellation (Anderson and Marchisio, 2012), and does not utilize any additional artificial augmentations. In addition to the contrastive task, GASSL also leverages the metadata available in RS imagery for a SSL task. This approach clusters together images by location and then predicts the latitude and longitude present in the metadata; the combination of this task coupled with contrastive learning across geographically aligned images results in a framework that attempts to explicitly learn time-agnostic representations of a given place. As with SeCo, the authors note that this approach learns temporally invariant features, but may struggle with tasks such as change detection (Ayush et al., 2021). Unlike the approaches below, latitude and longitude are not pre-processed in any way or treated as an additional modality.
4.2. Geographic Location as Augmentation
@@
Explicitly utilizing spatial data such as latitude and longitude provides value, but purely introducing them as features can result in models that perform poorly when asked to generalize to locations not within their training data (Klemmer et al., 2024; Jalayer et al., 2025). Therefore, a new generation of models dubbed “location encoders” have emerged, which once trained, take in latitude and longitude, then return an embedding of that location that quantifies and summarizes ground conditions of that location. Several models in this generation of location encoders rely on the CLIP framework (Radford et al., 2021) and treat ‘location’ as one modality, with a location-specific encoder to use in contrastive learning. This location encoder, in turn, distills the information extracted from the other mode, often an image obtained at that location.
@@
GeoCLIP (Cepeda et al., 2023) uses images taken at ground level from Flickr for one modality and spatial encodings for the other modality, then uses CLIP’s contrastive learning objective to learn effective representations of each modality. GeoCLIP uses Random Fourier Features (RFF), which have been shown to be highly effective in memorization and image reconstruction (Tancik et al., 2020), in order to store features learned from images into the location encoder. These Random Fourier Features allow GeoCLIP to learn high quality spatial encodings that capture information from the ground level photos.
@@
Contrastive Spatial Pretraining (CSP) (Mai et al., 2023b) also performs contrastive learning between ground level imagery and locations, but it explores a variety of techniques for generating positive and negative pairs that are explicitly tailored to spatially distributed data. In addition to CLIP’s standard in-batch negative sampling, CSP utilizes random negative location sampling, in which negative pairs for an image are generated by uniformly sampling locations at pre-training time; in this instance, the positive pair is still the corresponding location for the image. CSP also explores SimCSE-based sampling (Gao et al., 2022), in which two location encoders are initialized the same, but utilize different dropout masks. In this case, the same location is fed through each location encoder, and embeddings that point to the same initial location are considered positive pairs, while all other combinations are considered to be negative pairs. Each sampling approach forms a component of CSP’s contrastive loss function, resulting in a framework that learns nuanced representations of each modality.
@@
SatCLIP (Klemmer et al., 2024) also follows the CLIP approach while utilizing spatial encodings as one of its modalities and optical imagery captured from Sentinel-2 as the other. SatCLIP takes in all Sentinel-2 bands and does not limit itself purely to RGB data. SatCLIP does not utilize Random Fourier Features for the spatial encodings, and instead uses a SirenNet (Sitzmann et al., 2020) that relies on spherical harmonics (Rußwurm et al., 2023) to learn effective spatial encodings.
@@
Geo-Aligned Implicit Representations (GAIR) (Liu et al., 2025) takes this a step further by utilizing location as a way to align multispectral satellite images from Sentinel-2 and ground-level RGB images from the Global Streetscapes dataset (Hou et al., 2024). Ground-level imagery can provide a highly informative mode to train satellite imagery against, but a single satellite image’s scope is so large that hundreds or thousands of ground-level images can correspond to a single satellite image. In order to address this issue from a contrastive learning standpoint, GAIR learns Implicit Neural Representations (INR) (Xu et al., 2022) of a Sentinel-2 image. INR learns a continuous function for spatial coordinates to corresponding signals. These INR encodings of Sentinel-2 imagery are then contrasted against location encodings and ground level vision encodings. GAIR uses RFF (Tancik et al., 2020) for location encodings and a standard ViT for ground-level image encodings.
4.3. Different Sensors as Augmentation

### v2

@@ 3.3. Non-Image Data
Due to the ubiquity of the raster image format in remotely sensed data, current RS foundation models tend to be focused on using images to train, but a variety of other modalities do exist. Text, geographic location, and digital elevation maps (DEM) are all among potential additional modalities available in remote sensing at scale. It is conceivable that derivative information products such as population movement trajectories, social media data, and vector maps may be incorporated in enhancing foundation models. Any dataset with sufficient spatial coverage that contains geographic coordinates and a timestamp can feasibly be compared with corresponding satellite imagery. Such comparisons would further improve a model’s ability to interpret how satellite imagery pertains to ground level information. This spatiotemporal relationship results in a broad swathe of different modalities available for RS models to leverage. With an eye on these additional potential modalities, the rest of this manuscript focuses on the present and immediate future of foundation models in remote sensing.
@@ 4.1. Negative Sampling Models
One successful approach to SSL in CV is contrastive learning via negative sampling (Chen et al., 2020). Contrastive learning augments one image in two different ways, then a model is trained to recognize whether two images are augmentations of each other, i.e., are a positive pair. Any images that are not augmentations of one another are defined as negative pairs. For every batch, an image will have only one positive pair, and n-1 negative pairs, where n is the batch size. This approach can be thought of as a classification task where the model trains on classifying two augmentations as either a positive or negative pair. Within a given batch of samples, normalized-temperature cross entropy loss (InfoNCE Loss) (Chen et al., 2020) pulls together embeddings from the same original image and pushes apart embeddings that came from different images. This results in an encoder that learns sophisticated embeddings of images without the need for supervised learning. The encoder then can be ported to other downstream tasks, while reducing the need for labeled samples.
@@
SimCLR (Chen et al., 2020) achieved great success with this approach. SimCLR trained on ImageNet ILSVRC-2012 (Russakovsky et al., 2015) and introduced a non-linear projection head to improve the embedding quality of the layer before it. They also discovered that augmentations like color jitter and random cropping were particularly effective in forcing the model to learn embeddings that represented the data as a whole. Training without augmentations resulted in a model that learned dominant features, such as color distribution, and were able to trivialize the SSL task. Momentum Contrast (MoCo) (He et al., 2020) also utilizes contrastive learning, but maintains a dynamic memory bank or queue of negative samples, and updates it using a momentum encoder. This approach allows MoCo to effectively utilize smaller batch sizes than SimCLR, reducing memory requirements. Contrastive Language Image Pre-training (CLIP) (Radford et al., 2021) is a contrastive learning framework that utilizes image captions as a way to supervise learning image representations. CLIP trains two encoders, one for the text modality and one for the image modality. Given the multi-modal nature of this approach, contrastive learning to bring positive pairs together results in an alignment of the two modalities in the same embedding space. All of these models introduce concepts that have been leveraged by recent foundation models within the RS domain (Wang et al., 2022b).
4.2. Distillation Frameworks
@@
While contrastive learning via negative sampling achieves SOTA performance in CV, it generally requires large batch sizes. The number of negative pairs increases exponentially with relation to the batch size, and negative pairs are a key component of InfoNCE loss. This need for a large batch size results in significant hardware requirements, which is an obstacle to research for many teams. Distillation networks such as Bootstrap Your Own Latent (BYOL) (Grill et al., 2020) and Self-Distillation with No Labels (DINO) (Caron et al., 2021) were developed in part to address those hardware constraints.
Distillation simply refers to the concept of one model being trained on the outputs of another model (Phuong and Lampert, 2019). Distillation is not strictly an SSL method, as some research focuses on using distillation to ’compress’ knowledge from a large pre-trained model into a smaller model (Wu et al., 2024). Distillation leverages two encoders, dubbed the “teacher” and the “student”. When trying to ’compress’ logic, the teacher is some pre-trained model, but in the context of SSL, both student and teacher encoders are learning, albeit not at the same rate or with the same method, depending on the specific set up. Both are fed in augmented versions of an image to output embeddings. Then, contrastive loss is performed between the teacher and student embeddings and back propagated through the student network. The teacher network typically learns via the weighted average update system that was introduced in MoCo (He et al., 2020). Distillation models such as Joint-Embeddings Predictive Architecture (JEPA) (Assran et al., 2023) build on this framework by introducing a decoder after the student to predict the embeddings from the teacher. This approach leverages reconstruction loss in the latent space, but more distillation SSL approaches rely on contrasting embeddings from the student and teacher encoders (Grill et al., 2020; Chen et al., 2022; Jang et al., 2023; Oquab et al., 2024).
@@
The distillation methodology reduces the reliance on large batch sizes and extensive memory, making it more resource-friendly. It also outperforms comparable negative sampling models with fewer parameters (Grill et al., 2020). However, there is possibility of embedding collapse, as both the teacher and student networks learn based on the same loss (Chen and He, 2021). Embedding collapse describes a condition where a model generates embeddings that span a low-dimension subspace of the latent space, sometimes becoming so extreme as to only span a single point; this represents a significant loss of ability to capture information in embeddings (Jing et al., 2022). Frameworks introduce asymmetry through weighted average loss for the teacher (He et al., 2020) or an additional decoder for the student (Grill et al., 2020). This asymmetry between student and teacher is designed to prevent embedding collapse, but this phenomena is still an active area of research (Jain et al., 2022a).
4.3. Redundancy Reduction
Like the two contrastive approaches discussed above, redundancy reduction creates augmented views of an image and passes those views through an encoder to generate two embeddings, which it then attempts to align. But instead of simply aligning embeddings within the latent space, redundancy reduction generates an empirical cross correlation matrix between the two embeddings. The matrix is NxN, where N is the number of dimensions in the latent space, and each value ranges from -1 (perfect anti-correlation) to 1 (perfect correlation). Redundancy reduction loss is designed to make the resulting cross correlation matrix resemble the identity matrix.
Redundancy reduction approaches can differ in exactly how that loss is calculated. Barlow Twins (Zbontar et al., 2021) utilizes two terms: an invariance term and a redundancy reduction term. The invariance term forces the diagonal of the matrix to values of 1, while the redundancy reduction term forces the non-diagonal elements of the matrix to values of 0. This shares similar goals to InfoNCE loss, where the invariance term draws embeddings closer together (like positive pairs), and the redundancy reduction term forces embeddings further apart (like negative pairs). However, InfoNCE loss does this across a batch of embeddings, whereas Barlow Twins’ loss does so within a given embedding. VICReg (Bardes et al., 2022), builds on Barlow Twins’ intuition and includes three terms: a variance term to prevent collapse, an invariance term based on mean squared error between paired embeddings, and a covariance term to reduce redundancy among feature dimensions.
Redundancy reduction approaches, like distillation SSL, do not require as large of a batch size compared to negative sampling in order to learn effectively (Bardes et al., 2022). Barlow Twins also demonstrated an intriguing ability to more effectively leverage very-high-dimension embeddings: performance increased linearly with regard to the number of dimensions in the latent space. This trend was observed up to a latent dimensionality of 16K. (Zbontar et al., 2021).
4.4. Masked Image Modeling
Masked Image Modeling (MIM) refers to self-supervised frameworks that reconstructs masked portions of an input image. MIM is inspired by the Masked Language Modeling (MLM) paradigm introduced in BERT (Devlin et al., 2019). It is worth noting that this is different from autoregressive models such as GPT (Radford et al., [n. d.]), which learn to predict the next token in sequence rather than reconstruct masked inputs.
MIM generally entails the following steps:
Parts of a sample are masked.
The masked sample is run through an encoder to generate embeddings.
A decoder is asked to reconstruct the original sample from those embeddings.
Within MIM, several approaches have emerged beyond the widely used Masked Autoencoder (MAE) framework (He et al., 2021), including BEiT (Bao et al., 2022) (which predicts discrete visual tokens from a pre-trained tokenizer), SimMIM (Xie et al., 2022) (which reconstructs raw pixels of masked patches directly), and iBOT (Zhou et al., 2022) (which predicts masked patch embeddings through self-distillation). These differ primarily in their masking strategies, prediction targets, and encoder–decoder configurations. While the CV literature now spans this diversity of masked modeling techniques, MAE tends to be a popular option. Research in remote sensing has so far focused mainly on MAE-style methods due to their scalability and efficiency on large unlabeled satellite datasets (He et al., 2021), while incorporating ideas from other approaches to enhance representation learning. As such, we will tend to put central focus on MAE within the context of this paper.
@@ 4.4. Masked Image Modeling
MAE masks roughly 75% of the data to prevent trivial reconstruction from patch context (He et al., 2021). The generative nature of the task means that the batch size can be relatively small, making it resource-friendly. MIM approaches with established frameworks, such as Convolutional Neural Networks (CNNs) or Recursive Neural Networks (RNNs), replace a patch with a mask token, then feed in the whole image, including masked patches, into a network. But in MAE, the usage of ViTs allows for training pipelines to only feed encoders unmasked patches. Positional embeddings, in addition to patch embeddings, allows the ViT to avoid processing masked tokens. This shrinks the size of the necessary input layers and dramatically speeds up training time (He et al., 2021). Similarly, utilization of a lightweight decoder also forces the encoder to learn more accurate representations while speeding up runtime (He et al., 2021). These optimizations mean that the MAE approach scales remarkably well, resulting in models that can contain billions of parameters (Dehghani et al., 2023). Current CV research has identified a gap in the number of parameters between SOTA NLP and CV models. There are plans to improve the parameter count in CV foundation models even further (Dehghani et al., 2023).
@@ 5. Negative Sampling Models in Remote Sensing
While the underlying principles of negative sampling generally remain the same from CV to RS, RS images often share temporal or geospatial qualities. (Jain et al., 2022a). For example, Sentinel-1 and Sentinel-2 satellites can have images taken of the same location during the same day. These shared spatiotemporal attributes enables two potential advantages for RS foundation models:
Models can leverage imagery that comes from different RS sensors or different temporal acquisitions as naturally occurring augmentations.
It reduces the time required to do hyper-parameter tuning for augmentations.
The latter often adds significant overhead to training contrastive learning approaches (Jain et al., 2022b).
5.1. Temporal Pairing
@@
Pre-trained on RGB bands from Sentinel-2, Seasonal Contrast (SeCo) (Mañas et al., 2021) borrows heavily from SimCLR’s general framework and outperforms CV models such as MoCo on downstream land cover classification tasks. This indicates that there is implicit benefit in creating foundation models for the remote sensing domain as opposed to borrowing models directly from the CV field. Instead of only using hand-crafted image augmentations outlined in SimCLR (Chen et al., 2020), such as color jitter or gaussian blur, SeCo treats images of the same location at different points in time as natural augmentations of one another. However, SeCo does not simply rely on these natural augmentations. It also utilizes random cropping, random flipping, and color jitter to prevent the model from learning trivial representations of the data. The resulting model is essentially season- or time-agnostic, meaning it provides a general representation of data regardless of acquisition-time-of-year. This is helpful in seasonally invariant tasks, such as identifying land cover. However, it may not fit scenarios where retaining seasonal differences in embeddings for downstream tasks is essential, such as projecting forest fire recovery in an area.
@@
SeCo’s approach of contrastive learning between spatially aligned, temporally different RS image pairs has also been explored in Geographically Aware Self Supervised Learning (GASSL) (Ayush et al., 2021). However, GASSL trains on the Functional Map of the World (FMoW) (Christie et al., 2018) dataset, which contains multispectral imagery from the DigitalGlobe constellation (Anderson and Marchisio, 2012), and does not utilize any additional artificial augmentations. In addition to the contrastive task, GASSL also leverages the metadata available in RS imagery for a SSL task. This approach clusters together images by location and then predicts the latitude and longitude present in the metadata. The combination of this task coupled with contrastive learning across geographically aligned images results in a framework that attempts to explicitly learn time-agnostic representations of a given place. As with SeCo, the authors note that this approach learns temporally invariant features, and may struggle with tasks such as change detection (Ayush et al., 2021). Unlike the approaches below, latitude and longitude are not pre-processed in any way or treated as an additional modality.
5.2. Geolocation Pairing
@@
Explicitly utilizing spatial data such as latitude and longitude provides value, but purely introducing them as features can result in models that perform poorly when asked to generalize to locations not within their training data (Klemmer et al., 2024). Therefore, a new generation of models dubbed “location encoders” have emerged. Once trained, a location encoder can take in latitude and longitude and return an embedding of that location which quantifies and summarizes ground conditions present there. Several models in this generation of location encoders rely on the CLIP framework (Radford et al., 2021) and treat ‘location’ as one modality, with a location-specific encoder to use in contrastive learning. This location encoder, in turn, learns the information extracted from the other mode, often an image obtained at that location.
@@
GeoCLIP (Cepeda et al., 2023) uses images taken at ground level from Flickr for one modality and spatial encodings for the other modality. GeoCLIP then uses CLIP’s contrastive learning objective to learn effective representations of each modality. GeoCLIP uses Random Fourier Features (RFF), which have been shown to be highly effective in memorization and image reconstruction (Tancik et al., 2020), in order to store features learned from images into the location encoder. These Random Fourier Features allow GeoCLIP to learn high quality spatial encodings that capture information from the ground level photos.
@@
Contrastive Spatial Pretraining (CSP) (Mai et al., 2023b) also performs contrastive learning between ground level imagery and locations, but it explores a variety of techniques for generating positive and negative pairs that are explicitly tailored to spatially distributed data. In addition to CLIP’s standard in-batch negative sampling, CSP utilizes random negative location sampling, in which negative pairs for an image are generated by uniformly sampling locations at pre-training time. In this instance, the positive pair is still the corresponding location for the image. CSP also explores SimCSE-based sampling (Gao et al., 2022), in which two location encoders are initialized the same, but utilize different dropout masks. In this case, the same location is fed through each location encoder, and embeddings that point to the same initial location are considered positive pairs. All other combinations are considered to be negative pairs. Each sampling approach forms a component of CSP’s contrastive loss function, resulting in a framework that learns nuanced representations of each modality. CSP trains on ground-level imagery from the iNaturalist dataset (Van Horn et al., 2018).
@@
SatCLIP (Klemmer et al., 2024) also follows the CLIP approach while utilizing spatial encodings as one of its modalities and optical imagery captured from Sentinel-2 as the other. SatCLIP takes in all Sentinel-2 bands and does not limit itself purely to RGB data. SatCLIP does not utilize Random Fourier Features for the spatial encodings, and instead uses a SirenNet (Sitzmann et al., 2020) that relies on spherical harmonics (Rußwurm et al., 2024) to learn effective spatial encodings.
@@
Geo-Aligned Implicit Representations (GAIR) (Liu et al., 2025) takes this a step further by utilizing location as a way to align multispectral satellite images from Sentinel-2 and ground-level RGB images from the Global Streetscapes dataset (Hou et al., 2024). Ground-level imagery can provide a highly informative mode to train satellite imagery against. However, a single satellite image’s scope is so large that hundreds or thousands of ground-level images can correspond to a single satellite image. In order to address this issue from a contrastive learning standpoint, GAIR learns Implicit Neural Representations (INR) (Xu et al., 2022) of a Sentinel-2 image. INR learns a continuous function for spatial coordinates to corresponding signals. These INR encodings of Sentinel-2 imagery are then contrasted against location encodings and ground level vision encodings. GAIR uses RFF (Tancik et al., 2020) for location encodings and a standard ViT for ground-level image encodings.
5.3. Multisensor Pairing